OpenStudy (anonymous):

PLEASE HELP!!!!!!!!! WILL FAN AND GIVE MEDAL!!! Two students in your class, Wilson and Alexis, are disputing a function. Wilson says that for the function, between x = -1 and x = 1, the average rate of change is 0. Alexis says that for the function, between x = -1 and x = 1, the graph goes up through a turning point, and then back down. Explain how Wilson and Alexis can both be correct, using complete sentences

OpenStudy (loser66):

where is the problem from?

OpenStudy (anonymous):

A homework problem from algebra 2

OpenStudy (anonymous):

I have two word problems left and im stuck on both of them :/

OpenStudy (loser66):

since I don't see any problem on it, I don't know how to explain. Let see, if the function is a parabola y = -x^2 +1, then both them are correct. the average rate change of the function is defined by A =\(\dfrac{y_2-y_1}{x_2-x_1}\) but at x1 =-1, x2 = 1, y2=y1=0, so that the average rate change =0 although the graph is up and down then.

OpenStudy (anonymous):

Thats one of the things that confuses me the most. Im give very little and no problem and i dont really know what to do with what im given..

OpenStudy (loser66):

https://www.desmos.com/calculator/c9vb9bulpb

OpenStudy (anonymous):

@Loser66 can you help me with on more problem?